The following augmented matrix is in row echelon form and represents a linear system. Use back-substitution to solve the system if possible. 1 1-16 0 112 0 0 11 What is the solution to the linear system? Select the correct choice below and, if necessary, fill in the answer box to complete your choice. OA. The solution set is (Simplify your answer. Type an ordered triple.) There are infinitely many solutions. The solution set is x. Type an ordered triple. Type an expression using x as the variable.) O C. There is no solution
Answer #1

form augmented matrix,

z=1

.

y-z=2

y-1=2

y=3

.

x+y-z=6

x+3-1=6

x+2=6

x=4.

solution is (x,y,z)=(4,3,1)